I NEED HELP. PLEASE HELP ME SOLVE THIS. THANK YOU SO MUCH
9.54 Faced with rising fax costs, a firm issued a guideline that transmissions of 10 pages or more should be sent by 2-day mail instead. Exceptions are allowed, but they want the average to be 10 or below. The firm examined 75 randomly chosen fax transmissions during the next year, yielding a sample mean of 11.44 with a standard deviation of 2.15 pages. At the .05 level of significance, is the true mean greater than 10? Explain your decision.
Clearly state your decision on the null hypothesis.

Answer by stanbon(48568)

9.54 Faced with rising fax costs, a firm issued a guideline that transmissions of 10 pages or more should be sent by 2-day mail instead. Exceptions are allowed, but they want the average to be 10 or below. The firm examined 75 randomly chosen fax transmissions during the next year, yielding a sample mean of 11.44 with a standard deviation of 2.15 pages. At the .05 level of significance, is the true mean greater than 10? Explain your decision.
Clearly state your decision on the null hypothesis.
Ho: u <= 10
Ha: u > 10
--------------
test statistics: t(11.44) = (11.44-10)/[2.15/sqrt(75)] = 5.80
------------------
p-value = P(t > 5.80) = 0.0000000766
------------------
Since the p-value is less than 5%, reject Ho.
The true mean is greater than 10.
=====================================
Cheers,
Stan H.